HW1 Description The MATLAB script accompanying this assignment simulates the approximation of the periodic function f() defin please step by step
the main three question are the last three one thank you
Answer #1

(1) f(t) obtained by using fourier series of N-harmonics is

                                    f(t) = do + an cos(nwst) + bn sin(nwet) n=1 n=1

Here time period is

                                                                   T, = 21

So

                                                                       21 W = = = lrad/s

Highest frequency (in rad/s) is

                                                            mar = Nws = 200 rad)s

(2) For N = 0,

                                                                                   f(t) = ac

Here a0 is the DC value of the signal. From graph, we can see that average value of signal = 0. So we will get

                                                                           f(t) = 0

(3) New fnew(t) is

                                    N frey(t) = -10 + -an cos(nwgt) + -on sin(nwgt) n=1

We know from (2) that a0 = 0. Since f(t) is symmetric about t = 0, so it will not have any cos component. So

an = 0. So we get

                                                             にい arrows - (05

and

                                                           fnew(t) = -bn sin(nwgt) na

Or

                                             N fnew(t) = bn sin(-nwgt) n=1

Or

                                           N freu(t) = bn sin(nws(-+)) n=1

Or

                                                                 freu(t) = f(-1)

So new approximation will be time reversed form of original signal.